If a committee of 5 is to be chosen from 10 workers, what is the probability that it must include at least 2 women?
(There are 5 men and 5 women)

Answer by stanbon(75887)

If a committee of 5 is to be chosen from 10 workers, what is the probability that it must include at least 2 women?
(There are 5 men and 5 women)
----
# of ways to choose at least 2 women: 5C2 + 5C3 + 5C4 + 5C5 = 10+10+5+1 = 26
---------
# of random ways to select 5 randomly: 10C5 = 252
---------
P(at least 2 woman) = 26/252 = 13/126
--------------------
Cheers,
Stan H.
====================